Please Calculate the following and show work.

1- Calcuate the Angiomax doe for a patient with a normal CrCl, weight is 221 lb. Please calculate the doseage and volume for bolus as well as the drip rate?

2-Please explain how to mix Cardene for Intracoronary admin, for a mix of 100 mcg/ml. Stock is 25mg in 10 ml.
Answer #1

1)Given

Weight 221lb= 100.2(100kg)

Bolus dose: 0.75 mg/kg usually dose

so when 0.75mg=1kg

then?=100kg cross multiply and divide(100*0.75/1)

75mg of bolus dose will be administered

The vial contain 1ml=5mg

Then 75mg=?

15 ml to be given

To calculate drip rate

The standard calculation is 1.75mg/kg/hr

100*1.75/1=175 mg /hr will be transfused

Average time during PTCA is 4 hr

So =15*15/4*60=0.1 drop /minutes

2)Available 25mg/10ml

1mg=1000mcg (25mg=25000mcg)

So 100mcg will be 25000/100*10=25ml